Three point charges are located along the x axis as follows:  at Q1=30nC at (0,0,0), Q2=-27nC at...

  1. Three point charges are located along the x axis as follows:  at Q1=30nC at (0,0,0), Q2=-27nC at (12,0,0) and Q3 at (x,0,0) (all locations given on meters). Solve the following scenarios:
    1. Considering charge Q3=10nC , what is the total force applied to this charge if x=5 m.
    2. If the charge Q3=10nC , what must be the distance X so that the total force applied to the charge 3 is equal to Ft=-288.80nN .
  1. Using the three same charges from previous problem located at Q1(0,0,0), Q2(3, 5, -2) and Q3(1,2,3).  What is the total electric field at points:
    1. P(1,1,1)
    2. M(-2,4,-6)
Answer #1

0 - ܝܢ AR 47 Eo QR- FLD (30x19 toxicy i sam Due to point charge the electric force, Ē IR12 → unit vector Corresponding to ē Q- F2= Q3.02 (OR) 2) 45EU R = 1 R12 [ (5,0,0) -( 12,0,0)] Ř = -7a 11 = 1 (quxlc 9 (-29x109) F2 = (72 7an 7 4TEO 9x18x qoxi* Xnañ 3 F= 10X16?x 30X16*Pax 4TE (x2 [them an F = 9x1pMx 10x10 x 30x109 x2 - ^ 2700X16 an ILI 22 F = Q3.18 Q3.12 ă 4ā Eol R12 RFI= - 288.8x15 AM 4 Ro 2700 -2 ११.१४॥ २५30 (1-12)2 - -288.8 2700 2430 (A-12)2 M- 2100 + 288 - २५30 (x-12) २५30 भ+५५ - २५४ 2005) Electric freed due to point charge Q AR Ee 47Eo 1 R12 (A) E. at put,i,1) due to a(0,0,0) (30nc) 02 (3,5,-2) Q3 (1, 2, 3) E(6 9 Ez 9R Eas fQ 2 4Ats (| -23X1 x 9X107 & I- [1,1,1) [ 3,5,-2] 2 = -299 -49% +3 4 8| =439 ++ 16+9 R R - 23x9 E2 = [ 24 - 43ED = 90 E= 51.966 + 51.9699 +5190 -- 3-11 99-6-2243 (V5) 2 Es E= Et fit √5 [ aģ–2a3 ] 8.04 – 16.0893 V/m +4-6693 -- 8.04 ag -